Inspired Ltd manufactures spurs. Organisation policy requires factory overhead to be applied to the production of spurs using a predetermined rate based on budgeted direct labour hours. Budgeted cost of production (for 30000 units) for the year was, direct materials 225000, direct labour (6000 hours) 75000, fixed factory overhead 30000. Actual factory overhead incurred in the year was 72000. Actual direct labour hours was 6100. A) Calculate the factory overhead application rate (per direct labour hour) for the year. b) Calculate the total amount of factory overhead for the year applied to the production of spurs. c) Analyse under or over-applied overhead into two variances. Your answer must name the two variances, and indicate whether they are favourable /unfavourable
Answer #1

1) Solution: $11.50 per direct labor hour

Working: Fixed factory OH rate = (Fixed factory OH + Variable factory OH) / Budgeted direct labor hours

Fixed factory OH rate = (39,000 + 30,000) / 6,000

Fixed factory OH rate = 11.50 per direct labor hour

2) Total factory overhead for the year applied to the production of spurs:

Working: Fixed factory OH rate * Actual direct labour hours

= 11.50 per direct labor hour * 6100

= 70,150

3) Overhead underapplied: $1,850

Two variances:

Direct labor efficiency variance: $1,250 unfavorable

Variable OH efficiency variance: $500 unfavorable

Working: Overhead underapplied: Actual factory overhead - Applied factory overhead = 72,000 - 70,150 = $1,850

Direct labor efficiency variance: (6,000 - 6,100) * (75,000 / 6,000) = $1,250 unfavorable

Variable OH efficiency variance: (6,000 - 6,100) * (30,000 / 6,000) = $500 unfavorable